OpenStudy (anonymous):

Leon invested some money in an account paying 6% interest for one year. He invested 1500 more than this amount in an account paying 8%. How much did he invest in each account if the total interest earned in the year was 218?

OpenStudy (dumbcow):

Interest = .06x + .08(x+1500) = 218

OpenStudy (anonymous):

eh. it might be more intuitive to say. some money = x. 1.06x + (1500+x)*1.08 = (218+x+1500)